D Question 3 1 pts A stranded 103 kg man is motionless on a frozen, frictionless pond. To get to the shore. he throws his 8 kg backpack at 3 m/s away from the shore. How fast (speed in m/s) does the man move? ROUND ANSWER TO TWO DECIMAL PLACES.
Answer #1

Let the speed of man is v

Apply law of conservation of momentum

103 X 0 = 8×3 + 95 × v

24 = - 95v

v = - 24/95= - 0.25m/s